Percentage of Sales Models. Here are abbreviated financial statements for Planners Peanuts:

Income statement 2019

Sales

$2000

cost

$1500

Net Income

$500

Balance Sheet, year-end

2018

2019

2018

2019

Assets

$2500

$3000

Debt

$833

$1000

Equity

$1667

$2000

Total

$2500

Total

$2500

$3000

  1. If sales increase by 20% in 2020 and the company uses a strict percentage of sales planning model (meaning that all items on the income and balance sheet also increase by 20%, what must be the balancing item?

  2. What will be the value of this balancing item?

  3. What is the maximum possible growth rate for Planners Peanuts if the payout ratio remains at 50%? 3a. What if no Debt or equity is to be issued?

    ​​​​​​​3b. What if the firm maintains a fixed debt ratio, but issues no equity?
Answer #1

Answer-1:

The balancing item will be equity , its value = $2,000 + (20% of $2,000) = $2,000 + $400 = $2,400

Answer-3a:

Maximum possible growth rate if no Debt or equity is to be issued:

= retained earnings / assets

= ($500 - 50% of $500) / $3,000 = 83.33%

Answer-3b:

Maximum possible growth rate if the firm maintains a fixed debt ratio, but issues no equity:

= (retained earnings(1 +D/E)) / (debt + equity) = 6.37%